You are required to make a solution that contains 50 mmoles of buffer (its molecular mass...

  1. You are required to make a solution that contains 50 mmoles of buffer (its molecular mass is 121.1 g/mol), 300 mmoles of sodium chloride (its molecular mass is 58.44 g/mol), and 10% glycerol in a total volume of 0.75 L. Tell me (A) what weight of buffer and NaCl you will need, and (B) what volume of glycerol (diluted from a 100% stock concentration) you will need. Also, (C) what then makes up the additional volume to 0.75 L, and how much of that should you need (this assumes that the mass of buffer and NaCl make essentially no changes to the volume)?
Answer #1

A) The mass of the buffer and salt are calculated:

m buffer = n * MM = 0.05 mol * 121.1 g / mol = 6.055 g

m NaCl = 0.3 mol * 58.44 g / mol = 17,532 g

B) Glycerol volume is calculated:

V1 = C2 * V2 / C1 = 10% * 750 mL / 100% = 75 mL

C) It is completed with distilled water up to 750ml, the following are required:

V H2O = 750 - 75 = 675 mL

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.